OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations and provide the solution to the system. 5x - 9y = -16 2x + 6y = -16

OpenStudy (anonymous):

I would multiply the first equation by 2 and the second equation by -5. Once this is done, then add them, and the x's will cancel each other out. Then solve for y. Once y is found, sub that back into either of the equations and solve for x. 5x - 9y = -16 -->(2) 2x + 6y = -16 -->(-5) ---------------- 10x - 18y = - 32 (result of multiplying by 2) -10x - 30y = 80 (result of multiplying by -5) ----------------add - 48y = 48 y = -1 now sub -1 in for y in either of the original equations 5x - 9y = -16 5x - 9(-1) = -16 5x + 9 = -16 5x = -16 - 9 5x = - 25 x = -5 We can then check out answers by subbing in known valuesinto either equation 2x + 6y = -16 2(-5) + 6(-1) = -16 -10 -6 = -16 -16 = -16 (correct) SOLUTION : (-5,-1)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

no problem :)